OpenStudy (anonymous):

How do i simplify (8/7)/(2sqrt2) can someone show me the steps?

OpenStudy (anonymous):

\[\frac {\frac{8}{7}}{2\sqrt{2}}\rightarrow \frac{8}{7}* \frac{1}{2\sqrt{2}}\rightarrow \frac{8}{14\sqrt{2}}\] Now rationalize the denominator: \[\frac{8}{2\sqrt{2}}*\frac{\sqrt{2}}{\sqrt{2}}\rightarrow \frac{8 \sqrt{2}}{28}\rightarrow \frac {2\sqrt{2}}{7}\]

OpenStudy (anonymous):

sorry made a typo: \[\frac{8}{14\sqrt{2}}*\frac{\sqrt{2}}{\sqrt{2}}\]

OpenStudy (anonymous):

so the final answer is 2sqrt2/7?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

sure thing